He demanded and supply equetions for commodity Y are represented by
D=20-1/6p
S=18-1/8p
Required given that equlibrum price and quantity for commodity Y in a price market.
1
Expert's answer
2019-10-29T09:57:04-0400

Solution:

D=S


"20- \frac{1}{6} p=18- \frac{1}{8} p"

"20-18= \frac{1}{6} p - \frac{1}{8} p"

"\frac{1}{24} p=2"

P=48.


"Q=20- \frac{1}{6} \times48"

"Q=12"

Answer: price 48 and quantity 12.